A small pith ball has a charge of 0.001 C. What is the electric field 0.3 m from the ball? that k-8.988x10A9 N m 2/CA2 Remember 3.0x107 N/C a. 1.0x1048 N/C b. 0.1 N/C C. It depends on the charge of the object placed in the field. d.
Answer #1

Charge of the pith ball, q =0.001 C

distance from the ball, r = 0.3 m

Electric field 0.3m from the ball, E = k ( q2 / r2 ) = 8.988*10^(9) * (((0.001)^2) / ((0.3)^2)) = 105 N/C (ans)

None of the given options are correct.